# Neural network prediction of model parameters for strong lensing samples from Hyper Suprime-Cam Survey

### v2

@@ 4 Results and Discussion
We train our network using simulated lenses and their true SIE lens model parameters, namely, the Einstein radius, axis ratio and position angle (see Appendix A for conventions). We further convert the axis ratio and the position angle into the two components of the ellipticity ($e_{x}$ and $e_{y}$) as discussed in the Appendix A. Amongst the three parameters Einstein radius is the most robustly predicted parameter as it depends on the radial angular separations of the lensed arcs from the central lens galaxy. However, the axis ratio and position angle are related to lensed arc properties such as their configuration, their orientation and their multiplicity and thus, are difficult to infer accurately. Presence of the lens light may further worsen their prediction.
We train our network with four different simulated samples with increasing realism. The PureSims sample, where we convolve our simulated arcs with the PSF and add Poisson noise to them without adding any background noise, is a very ideal case. In this case we expect the network to be thrown off due to the presence of background galaxies, noise fluctuations in real images or imperfect subtraction of the lens light. In order to address these issues incrementally, we have used additional training samples. In the GauNoise training sample, we generate Gaussian background noise equivalent to our expectation in the Subaru HSC survey and add it to the PureSims sample, such that the network learns about the presence of noise and becomes insensitive to noise peaks. The HSCempty training sample, comprises of lensed source images added to cutouts from real HSC data which do not contain bright objects near the centre, and thus contains both noise and unassociated galaxies in the cutout. Finally, we also create a LensLight training sample, where the simulated lensed sources are added to the HSC image of the galaxy which was used to model the gravitational lensed images, in order to test whether lens light can further provide useful information for parameter estimation.
@@
The performance of our network on PureSims, GauNoise, HSCempty and LensLight test samples can be compared using the statistical quantities defined above and the standard deviation ($\sigma$), bias ($b$) and outlier fraction ($f_{\rm{out}}$) are shown in the three different rows of Fig. 6, respectively. On the other hand, the three columns correspond to the parameters $\theta_{\rm E}$, $e_{x}$ and $e_{y}$, respectively.
@@
The performance on the real SuGOHI lenses of our network when trained on the different samples is shown in the bottom panels of Figs. 2 -  5. Even though the network trained on PureSims gave the least scatter on the simulated test sample, it fares very poorly on the real SuGOHI lenses, likely due to the missing features in the training such as background noise and the presence of other galaxies. In particular, the network prediction for the Einstein radius shows no correlation with the Einstein radius of the real lenses. We believe that this is likely a result of the network latching on to either noise or other unassociated galaxies in the image as being a part of the lensed images of the source or due to imperfect subtraction of the lens light from the real SuGOHI lenses.
The network trained on the additional training samples help address these issues one by one. The performance on the real SuGOHI lenses of the network trained on the GauNoise sample is shown in the bottom panels of Fig. 3. Using a training sample which includes Gaussian noise improves the performance of the network on the real SuGOHI test images, where a positive correlation emerges in the network predictions of the Einstein radius compared to those from YL .
The performance on the real SuGOHI lenses of the network trained on the Lenslight sample is shown in the bottom panel of Fig. 5. When compared to the results from the GauNoise sample, we see lesser correlation in the prediction of the Einstein radius on the real SuGOHI sample, and an almost flat prediction of the ellipticity components. Although this sample resembles well with the realistic scenario, as it has the lens light, the noise as well as other unassociated galaxies, it appears that the network is unable to adequately separate the light coming from the central lens galaxy, which either outshines or contaminates the background lensed arcs due to blending. This shows the importance of accurate modelling and subtraction of the lens light before the images are fed to the network.
The results for the network trained using the HSCempty sample, which has all the desirable features like HSC foregrounds and backgrounds but no lens light to contaminate the arcs, on the real SuGOHI lenses are shown in the bottom panel of Fig. 4. We present the corresponding lens model parameters estimated by the network and YL , along with their errors, in the Table 2 (see Appendix D). Training on such samples improves the results for the SuGOHI test sample than when trained on the Lenslight sample and a performance comparable to the GauNoise sample.
@@ 4.1 Network performance comparison of 10 real SuGOHI lenses commonly modelled by S19 and S23b
    HSCJ015618-010747 : In this case, the models of YL , S19 and S23b (see their fig. B.1) are visually similar. The $e_{x}$ and $e_{y}$ values predicted by S23b, YL and network are within the quoted errorbars, while the predictions for $\theta_{\rm E}$ are close enough for all the methods. The ellipticity components predicted by S19 are close to zero and deviate from other results.
@@
    HSCJ020241-064611 : For this lens, the models of YL , S19 and S23b (see their fig. B.3) look qualitatively similar with one image in the north and the other in the south direction. The values of $\theta_{\rm E}$ for all four methods, are consistent within the quoted errorbars due to the similarity of models containing two compact images, whereas the ellipticity components are not constrained well enough .
@@
    HSCJ142720+001916 : For this lens, S19 and S23b (see their fig. B.22) models appear more accurate with an extended source whereas YL model is more circular with a compact source. The models from all methods agree well on $\theta_{\rm E}$ and the predictions for $e_{x}$ are also within the errorbar from our network, while the parameter $e_{y}$ is quite poorly constrained.
@@
    HSCJ230335+003703 : For this quad lens system, S19 model looks quite accurate with distinct four images of a compact source. The lens model of YL does not seem accurate. It has an extended source and the model looks more circular than the actual configuration of the lensed images. The model of S23b (see their fig. B.28) also does not do justice to the actual configuration. The $\theta_{\rm E}$ for all the methods is roughly consistent within the quoted errorbars. The inferred $e_{y}$ from YL , S19 and S23b agree well but our network deviates from this, whereas for $e_{x}$ the predictions from YL and S19 are within the errorbar from our network.
After comparing results from the three conventional modelling methods, namely, YL , S19 and S23b with our network, we reach a similar conclusion to S23b. Even if sophisticated lens modelling methods are used on the same SuGOHI lenses with identical data quality, barring the Einstein radius, there is little consensus amongst the inferred parameters. The parameters of the ellipticity and its position angle, are likely heavily prior and or algorithm dependent. For instance, insufficiencies in modelling the light profile of the lens galaxy and the source, the choice of including or excluding the external shear and using different combinations of the broad-band data, can all heavily affect the answers beyond the quoted statistical uncertainties. In addition, the actual mass model of the galaxy can be quite complicated than the SIE+external shear model. However, in the absence of the ground truth, one cannot assess which of the methods and their results are more accurate. It may well be that the limitation is inherent to the quality and resolution of the ground-based survey data and better accuracy on the parameters is not possible unless working with data from space based surveys like the Hubble Space Telescope (e.g., Hezaveh et al., 2017).
@@ 4.2 Comparison of our network and results with studies from the literature
Our work has numerous similarities but also important differences to the growing literature where a neural network is used to analyse parameters of strong lensing systems, particularly from imaging surveys resembling the HSC survey (e.g., S21, S23a, S23b). We present a brief comparison of these methods in this section. The analysis done in P19 focuses on ground based imaging from the Rubin LSST survey, and given that the survey is yet to start, rely on training and testing their network on images where both the lens and the source images are simulated. The studies of S21, S23a, simulate source galaxies and inject them around real lens galaxies from the Subaru HSC survey. The study of S23b applies the network trained and developed in S23a to real SuGOHI lenses.
@@
In S21, S23a and S23b, they simulate and analyse HSC-like images, similar to our analysis which enables a fairer comparison with these studies than P19. The analysis carried out in S21 using a simple CNN (see fig. 6 in S21) is quite similar to our analysis using the LensLight sample (see the upper row of Fig. 5). Comparing these two plots, we infer that the predictions for the components of ellipticities from both the studies appear qualitatively similar. The use of lens light to determine the ellipticity works best as long as there is little scatter between the ellipticity of the lens light and the lens potential. For the Einstein radius, although S21 results appear slightly better than our results with the LensLight sample, they are comparable with our predictions for the Einstein radius with the HSCempty sample (see the upper row of Fig. 4).
In S23a, the authors build upon the results of S21 with a ResNet which improves the recovery of any spatial offsets between the lens light and mass distribution. However, although the ResNet improved their results on the $x$ and $y$ dramatically (see fig. 4 in S23a) compared to S21, the results on the components of ellipticity remain comparable and also quantitatively similar to our network predictions with the LensLight sample (see the upper row of Fig. 5). The results in S23a, corresponding to the Einstein radius (see fig. 4 in S23a) seem to slightly degrade compared to S21 and also seem to match the performance of the inference of the Einstein radius from our network trained on the LensLight sample. Although S23a additionally include external s, hhe network has difficulty in the inference of the external shear. The performance of the network from S23a with traditional lens modelling methods on grade A SuGOHI sample of lenses was presented in S23b and can be compared with the comparison between our network and traditional lens modelling method (see Appendix  C). We see that the networks in both these studies have comparable performance in reproducing the results from traditional modeling methods for both the Einstein radius and the two components of the ellipticity.
@@ 5 Summary and Conclusion
We developed a simple CNN to analyse the strong lenses from the HSC data, a precursor to Rubin LSST, and estimate the parameters of the SIE lens mass model, namely, the Einstein radius, the axis ratio and the position angle of the major axis of the mass distribution. We first trained and tested our network on 60000 HSC-like galaxy-scale simulated lenses. We prepared four different training samples with increasing degree of realism: PureSims sample included only the arcs with Poisson noise, the GauNoise sample further included background Gaussian noise, the HSCempty sample included the lensed features on top of randomly selected cutouts from HSC with central “empty" regions, and the LensLight sample also included the light from the lens galaxy.
@@ Acknowledgements
We thank Stefan Schuldt, Francisco Villaescusa-Navarro, Vibhore Negi, Shreejit Jadhav, Vishal Upendran and Navin Chaurasiya along with Sukanta Bose for useful discussions on the project. We also thank Stefan Schuldt for sharing the data. PG acknowledges financial support provided by the University Grants Commission (UGC) of India. We acknowledge the use of the high performance computing facility Pegasus at IUCAA for this work. NY and AK thank financial support by Japan Science and Technology Agency AIP Acceleration Research Grant Number JP20317829.
@@ Appendix C Performance of our LensLight sample on the grade A SuGOHI lenses
We show the performance of our LensLight sample on 182 SuGOHI (grade A and B) lenses in the lower panel of the Fig. 5. In S23b, they train their ResNet on a sample qualitatively similar to our LensLight sample and test it on a sample of 31 grade A SuGOHI lenses. In order to compare with S23b, here we show the performance of our LensLight sample on a sample 25 grade A suGOHI lenses. Our grade A sample is smaller than the S23b sample, since YL does not identify a few of these 31 grade A lenses as lenses and we do not have the parameters from YL to compare our results with for those lenses. If we compare the results shown in the Fig. 11 (we compare our network predictions with YL ) with the fig.4 in S23b (they compare their network predictions with GLEE & GLAD), we can see that our results are comparable with theirs for the Einstein radius and the components of the ellipticity.
@@
Appendix D Lens model parameters for SuGOHI lenses

### v3

@@ 4 Results and Discussion
We train our network using simulated lenses and their true SIE lens model parameters, namely, the Einstein radius, axis ratio and position angle (see Appendix A for conventions). We further convert the axis ratio and the position angle into the two components of the ellipticity ($e_{x}$ and $e_{y}$) as discussed in the Appendix A. Amongst the three parameters Einstein radius is the most robustly predicted parameter as it depends on the radial angular separations of the lensed arcs from the central lens galaxy. However, the axis ratio and position angle are difficult to infer accurately. The presence of lens light may further worsen their prediction.
We train our network with four different simulated samples with increasing realism. The PureSims sample, where we convolve our simulated arcs with the PSF and add Poisson noise to them without adding any background noise, is a very ideal case. In this case we expect the network to be thrown off due to the presence of background galaxies, noise fluctuations in real images or imperfect subtraction of the lens light. In order to address these issues incrementally, we have used additional training samples. In the GauNoise training sample, we generate Gaussian background noise equivalent to our expectation in the HSC survey and add it to the PureSims sample, such that the network learns about the presence of noise and becomes insensitive to noise peaks. The HSCempty training sample, comprises of lensed source images added to cutouts from real HSC data which do not contain bright objects near the centre, and thus contains both noise and unassociated galaxies in the cutout. Finally, we also create a LensLight training sample, where the simulated lensed sources are added to the HSC image of the galaxy which was used to model the gravitational lensed images, in order to test whether lens light can further provide useful information for parameter estimation.
@@
The performance of our network on PureSims, GauNoise, HSCempty and LensLight test samples can be compared using the statistical quantities defined above and the standard deviation ($\sigma$), bias ($b$) and outlier fraction ($f_{\rm{out}}$) are shown in the three different rows of Fig. 7, respectively. The three columns correspond to the parameters $\theta_{\rm E}$, $e_{x}$ and $e_{y}$, respectively.
@@
The performance on the real SuGOHI lenses of our network when trained on the different samples is shown in the bottom panels of Figs. 3 -  6. Even though the network trained on PureSims gave the least scatter on the simulated test sample, it fares very poorly on the real SuGOHI lenses, likely due to the missing features in the training such as background noise and the presence of other galaxies. In particular, the network prediction for the Einstein radius shows no correlation with the Einstein radius of the real lenses. We believe that this is likely a result of the network latching on to either noise or other unassociated galaxies in the image as being a part of the lensed images of the source or due to imperfect subtraction of the lens light from the real SuGOHI lenses.
The networks trained on more complex training samples help address these issues one by one. The performance on the real SuGOHI lenses of the network trained on the GauNoise sample is shown in the bottom panels of Fig. 4. Using a training sample which includes Gaussian noise improves the performance of the network on the real SuGOHI test images, where a positive correlation emerges in the network predictions of the Einstein radius compared to those from YL .
The performance on the real SuGOHI lenses, for a network trained on the Lenslight sample, is shown in the bottom panel of Fig. 6. When compared to the results from the GauNoise sample, we see lesser correlation in the prediction of the Einstein radius on the real SuGOHI sample, and an almost flat prediction of the ellipticity components. Although this sample resembles well with the realistic scenario, as it has the lens light, the noise as well as other unassociated galaxies, it appears that the network is unable to adequately separate the light coming from the central lens galaxy, which either outshines or contaminates the background lensed arcs due to blending. This shows the importance of accurate modelling and subtraction of the lens light before the images are fed to the network.
The results for the real SuGOHI lenses with the network trained using the HSCempty sample, which has all the desirable features like HSC foregrounds and backgrounds but no lens light to contaminate the arcs, are shown in the bottom panel of Fig. 5. We present the corresponding lens model parameters predicted by the network and YL , along with their errors, in the Table 2 (see Appendix D for details). Training on such samples improves the results for the SuGOHI test sample than when trained on the Lenslight sample and a performance comparable to the GauNoise sample.
@@ 4.1 Network performance comparison of 10 real SuGOHI lenses commonly modelled by S19 and S23b
    HSCJ015618-010747 : In this case, the models of YL , S19 and S23b (see their fig. B.1) are visually similar. The $e_{x}$ and $e_{y}$ values predicted by S23b, YL and network are within the quoted errorbars, while the predictions for $\theta_{\rm E}$ are similar for all the methods. The ellipticity components predicted by S19 are close to zero and deviate from other results.
@@
    HSCJ020241-064611 : For this lens, the models of YL , S19 and S23b (see their fig. B.3) look qualitatively similar with one image in the north and the other in the south direction. The values of $\theta_{\rm E}$ for all four methods, are consistent within the quoted errorbars due to the similarity of models containing two compact images, whereas the ellipticity components are not.
@@
    HSCJ142720+001916 : For this lens, S19 and S23b (see their fig. B.22) models appear more accurate with an extended source whereas YL model is more circular with a compact source. The models from all methods agree well on $\theta_{\rm E}$ and the predictions for $e_{x}$ are also within the errorbars, while the parameter $e_{y}$ is quite poorly constrained.
@@
    HSCJ230335+003703 : For this quad lens system, S19 model looks quite accurate with distinct four images of a compact source. The lens model of YL does not seem accurate. It has an extended source and the model looks more circular than the actual configuration of the lensed images. The model of S23b (see their fig. B.28) also does not do justice to the actual configuration. The $\theta_{\rm E}$ for all the methods is consistent within the quoted errorbars. The inferred $e_{y}$ from YL , S19 and S23b agree well but our network deviates from this, whereas for $e_{x}$ the predictions from YL and S19 are within the errorbars.
After comparing results from the three conventional modelling methods, namely, YL , S19 and S23b with our network, we reach a similar conclusion to S23b. Even if sophisticated lens modelling methods are used on the same SuGOHI lenses with identical data quality, barring the Einstein radius, there is little consensus amongst the inferred parameters. The parameters of the ellipticity and its position angle, are likely heavily prior and or algorithm dependent. For instance, insufficiencies in modelling the light profile of the lens galaxy and the source, the choice of including or excluding the external shear and using different combinations of the broad-band data, can all heavily affect the answers beyond the quoted statistical uncertainties. In addition, the actual mass model of the galaxy can be more complicated than the SIE+external shear model. However, in the absence of the ground truth, one cannot assess which of the methods and their results are more accurate. It may well be that the limitation is inherent to the quality and resolution of the ground-based survey data and better accuracy on the parameters is not possible unless working with data from space based surveys like the Hubble Space Telescope (e.g., Hezaveh et al., 2017).
@@ 4.2 Comparison of our network and results with studies from the literature
Our work has numerous similarities but also important differences to the growing literature where a neural network is used to analyse parameters of strong lensing systems, particularly from imaging surveys resembling the HSC survey (e.g., S21, S23a, S23b). We present a brief comparison of these methods in this section. The analysis done in P19 focuses on ground based imaging from the LSST survey, and given that the survey is yet to start, rely on training and testing their network on images where both the lens and the source images are simulated. The studies of S21, S23a, simulate source galaxies and inject them around real galaxies from the HSC survey. The study of S23b applies the network trained and developed in S23a to real SuGOHI lenses.
@@
In S21, S23a and S23b, they simulate and analyse HSC-like images, similar to our analysis which enables a fairer comparison with these studies than P19. The analysis carried out in S21 using a simple CNN (see fig. 6 in S21) is quite similar to our analysis using the LensLight sample (see the upper row of Fig. 6). Comparing these two plots, we infer that the predictions for the components of ellipticities from both the studies appear qualitatively similar. The use of lens light to determine the ellipticity works best as long as there is little scatter between the ellipticity of the lens light and the lens potential. For the Einstein radius, although S21 results appear slightly better than our results with the LensLight sample, they are comparable with our predictions for the Einstein radius with the HSCempty sample.
In S23a, the authors build upon the results of S21 with a ResNet which improves the recovery of any spatial offsets between the lens light and mass distribution. However, although the ResNet improved their results on the $x$ and $y$ dramatically (see fig. 4 in S23a) compared to S21, the results on the components of ellipticity remain comparable. The results in S23a, corresponding to the Einstein radius (see fig. 4 in S23a) seem to slightly degrade compared to S21. Although S23a additionally include external shear, the network has difficulty in the inference of the external shear. We find that the performance of our network is better or comparable to the case of S23a. The performance of the network from S23a with traditional lens modelling methods on grade A SuGOHI sample of lenses was presented in S23b and can be compared with the comparison between our network and traditional lens modelling method (see Appendix  C). We see that the networks in both these studies have comparable performance in reproducing the results from traditional modeling methods for both the Einstein radius and the two components of the ellipticity.
@@ 5 Summary and Conclusion
We developed a simple CNN to analyse the strong lenses from the HSC data, a precursor to LSST, and estimate the parameters of the SIE lens mass model, namely, the Einstein radius, the axis ratio and the position angle of the major axis of the mass distribution. We first trained and tested our network on 60000 HSC-like galaxy-scale simulated lenses. We prepared four different training samples with increasing degree of realism: PureSims sample included only the arcs with Poisson noise, the GauNoise sample further included background Gaussian noise, the HSCempty sample included the lensed features on top of randomly selected cutouts from HSC with central “empty" regions, and the LensLight sample also included the light from the lens galaxy.
@@ Acknowledgements
We thank Stefan Schuldt, Francisco Villaescusa-Navarro, Vibhore Negi, Shreejit Jadhav, Vishal Upendran and Navin Chaurasiya along with Sukanta Bose for useful discussions on the project. We also thank Stefan Schuldt for sharing the data. PG acknowledges financial support provided by the University Grants Commission (UGC) of India. We acknowledge the use of the high performance computing facility Pegasus at IUCAA for this work. NY and AK thank financial support by Japan Science and Technology Agency AIP Acceleration Research Grant Number JP20317829 and JSPS Kakenhi 24H00221.
@@ Appendix C Performance of our LensLight sample on the grade A SuGOHI lenses
We show the performance of our LensLight sample on 182 SuGOHI (grade A and B) lenses in the lower panel of the Fig. 6. In S23b, they train their ResNet on a sample qualitatively similar to our LensLight sample and test it on a sample of 31 grade A SuGOHI lenses. In order to compare with S23b, here we show the performance of our LensLight sample on a sample 25 grade A suGOHI lenses (see Fig. 12). Our grade A sample is smaller than the S23b sample, since YL does not identify a few of these 31 grade A lenses as lenses and we do not have the parameters from YL to compare our results with for those lenses. If we compare the results shown in the Fig. 12 (we compare our network predictions with YL ) with the fig.4 in S23b (they compare their network predictions with GLEE & GLAD), we can see that our results are comparable with theirs for the Einstein radius and the components of the ellipticity.
@@
Appendix D Lens model parameters for SuGOHI lenses and Comparison with YL
@@
To evaluate the reliability of our network predictions on real data, we evaluate its performance on a test sample of 182 SuGOHI lenses (grade A + B), processed using YL for central lens light subtraction and mass modeling. Based on visual inspection of the YL output, we classify the lenses into two categories: GOOD, comprising 122 lenses where both the lens light subtraction and mass modelling appeared to perform well; and POOR, comprising 60 lenses where one or both steps appeared suboptimal as shown in the upper panel of Fig. 13.
To explore whether the networks predictions correlate differently with the YL estimates across these categories, we calculate the perpendicular distances from the $x=y$ line in parameter comparison plots and analyse the normalised distributions of their logarithms for GOOD and POOR lenses. A Kolmogorov–Smirnov (KS) test reveals statistically significant differences between the GOOD and POOR distributions, with $p$-values of 0.017 for $\theta_{\rm E}$, 0.003 for $e_{x}$, and 0.014 for $e_{y}$ (see lower panel of Fig. 13). These results indicate that the network predictions align more closely with YL estimates for the lenses where the lens light subtraction and mass modelling by YL is visually assessed to be more reliable.
The average time required for YL to perform a quick lens light subtraction is approximately 15 s per lens, with more refined procedures taking even longer. Given the large volume of lenses expected from upcoming surveys such as LSST, this presents a computational challenge. To address this, there is a need to develop faster lens light subtraction techniques, potentially leveraging machine learning approaches such as convolutional autoencoders like U-Net architectures, to enable scalable preprocessing at survey scale. Developing such an algorithm would require substantial effort and is therefore beyond the scope of this work. However, it represents a promising direction for future research.
@@
Figure 13: Upper panel: Performance of our network trained on the HSCempty sample, evaluated on a test set of 182 SuGOHI (grade A+B) lenses. These lenses were processed using YL for central lens light subtraction. Based on visual inspection of the output of YL, we classify the lenses into two categories, GOOD: lenses where YL performs both lens light subtraction and modelling successfully (122 lenses, shown in black) and POOR: lenses where either component fails (60 lenses, shown in cyan).
Lower panel: Normalised cumulative distributions of the perpendicular distances from the $x=y$ line for the three lens parameters for the GOOD and POOR samples. We perform a Kolmogorov–Smirnov (KS) test to assess whether the two distributions differ significantly, and report the corresponding $p$-value.